### v1

# Interactive Program Synthesis for Modeling Collaborative Physical Activities from Narrated Demonstrations

Abstract.
Teaching systems physical tasks is a long-standing goal in HCI, yet most prior work has focused on non-collaborative physical activities. Collaborative tasks introduce added complexity, requiring systems to infer users’ assumptions about their teammates’ intent—an inherently ambiguous and dynamic process. This necessitates representations that are interpretable and correctable, enabling users to inspect and refine system behavior. We address this challenge by framing collaborative task learning as a program synthesis problem. Our system represents behavior as editable programs and uses narrated demonstrations—paired physical actions and natural language—as a unified modality for teaching, inspecting, and correcting system logic without requiring users to see or write code. The same modality is used for the system to communicate its learning to users. In a within-subjects study, 20 users taught multiplayer soccer tactics to our system. 70% (14/20) of participants successfully refined learned programs to match their intent and 90% found it easy to correct the programs. The study surfaced unique challenges in representing learning as programs and in enabling users to teach collaborative physical activities; we discuss these issues and outline mitigation strategies.
program synthesis, interactive programming, probabilistic programming language, machine learning, artificial intelligence, human computer interactions
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; .; .isbn: 978-1-4503-XXXX-X/2018/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
Figure 1. Our system learns collaborative physical activities from narrated mixed reality demonstrations, synthesizes programs with domain-specific APIs, and demonstrates back to the user for feedback and refinement.:
1. Introduction
Enabling users to teach systems how to perform physical tasks has been a long-standing goal in human-computer interaction (HCI). While significant progress has been made in teaching individual skills, less attention has been given to enabling non-technical users to teach systems complex, collaborative activities, e.g., in sports, construction, and manufacturing. In this paper, we introduce a system that addresses several key challenges in this area. We aim to: (1) develop a system that can effectively acquire and model these collaborative activities in the physical world; and (2) enable users with no technical background to interactively teach and correct the system’s learning to align its behavior with their intent. Addressing these problems would allow for the creation of systems that can scalably teach, assist, and guide people in complex, real-world collaborative scenarios in the future.
However, realizing such a system presents significant technical challenges. First, learning collaborative activities entails substantially greater complexity compared to single-user tasks.. In HCI, systems for learning physical tasks have predominantly relied on demonstrations and, more recently, natural language. While teaching via demonstrations can capture the nuances of physical movement, it often fails to infer the underlying logic or user intent (Cypher, 1993). Conversely, natural language can articulate high-level goals but struggle to ground abstract commands into specific, continuous trajectories in the physical world—a classic instance of the symbol grounding problem (Harnad, 1990; Bisk et al., 2020). Collaboration further increases this complexity since user’s assumptions on collaborators’ intent also need to be accounted for. Also, the constraint that the system must be teachable by users without a technical background necessitates that its internal models of physical activities be both interpretable and directly editable through an interface.
The existing literature on systems for representing collaborative physical activities is sparse. The most closely related works, such as GhostAR (Cao et al., 2019) and VisCourt (Cheng et al., 2024), leverage mixed reality (MR) to generate virtual avatars for demonstrating interactions. This approach provides the system with a fully observable view of the behaviors of, and interactions between, the user and their collaborators. However, rather than teaching collaborative tasks, users configure the interactions to generate tutorials. For example, in VisCourt (Cheng et al., 2024), a basketball coach selects and parameterizes pre-defined tactical plays from a menu-based interface. In our problem of interest, we aim to have users teach collaborative activities to the system using natural modalities (e.g., language and demonstrations).
We present an interactive system that enables non-programmers to model collaborative physical activities by synthesizing programs from narrated demonstrations. Our choice of this input modality—physical demonstrations paired with concurrent natural language—is informed by prior work showing that it facilitates more accurate and robust learning (Tung et al., 2018; Waytowich et al., 2019; Yu and Mooney, 2023); narration disambiguates the intent behind an action, while the demonstration grounds language in the physical world. As summarized in Fig. 1, our system treats narrated demonstration as a key interaction primitive across the entire teaching loop—for initial instruction, inspection, and correction—enabling users to align the program’s behavior with their intent without interacting with raw source code. The system allows a user to first teach a task via a narrated demonstration in mixed reality (MR). The system processes this recording with a large vision-language model (LVLM) to generate an initial program. The system communicates its learning via decision flow diagram (visualizing the underlying code structure) and program execution which provides narrated demonstrations to the user. To correct the program, the user provides feedback by providing narration along with physical demonstration via annotations.
We evaluated our system in a within-subjects study with 20 non-programmer participants who used it to teach multi-player soccer tactics in a mixed reality environment. Our results show that users, leveraging the iterative feedback loop, were able to successfully teach and significantly refine the system’s understanding of complex, collaborative behaviors to align with their intent. Users also reported ease of use in teaching and correcting the system. This paper makes the following key contributions:
  1. (1)
    An interactive programming synthesis system for modeling collaborative physical activities that enables:
    1. (a)
      non-programmers to teach and correct the system via narrated demonstrations, and
    2. (b)
      the system to effectively communicate its learning to users via narrated demonstrations.
  2. (2)
    A user study demonstrating that non-programmers can use the proposed interactive framework to:
    1. (a)
      easily teach, inspect, and correct the system so as to model complex collaborative physical activities as programs, and
    2. (b)
      align program behaviors with the intent of users
2. Related Work
2.1. Tutorial Authoring Systems
Augmented Reality (AR) has long been envisioned for delivering tutorials and task guidance. Seminal systems demonstrated knowledge-based overlays for step guidance (Feiner et al., 1993) and proactive, sensor-driven instructions for physical assembly (Antifakos et al., 2002). However, these early systems focused on the delivery of guidance, with authoring relegated to experts rather than end users. A significant thread in HCI research has focused on empowering practitioners to author their own tutorials. For example, InstruMentAR (Liu et al., 2023b) and AdapTutAR (Huang et al., 2021) enable experts to record embodied demonstrations to produce step-wise AR overlays for machine operation, with AdapTutAR further tailoring guidance to the learner’s state. Beyond AR/VR, the community has explored auto-authoring from demonstrations in mixed media. TutoriVR (Kumaravel et al., 2019) embeds 2D videos within VR workspaces accompanied by 3D contextual cues, while VAAR (Yamaguchi et al., 2020) reconstructs a 3D assembly graph from 2D video to align AR animations. TutorialLens (Kong et al., 2021) advances this by directly authoring AR tutorials from synchronized narration and demonstration, capturing finger trajectories and speech to structure the steps.
A critical distinction in recent work is the role of natural language. While much prior work simply records and replays narration to augment steps, newer systems analyze language to infer procedural structure. For instance, Truong et al. automatically derive two-level hierarchical procedures from instructional videos (Truong et al., 2021), and TutoAI (Chen et al., 2024) systematizes pipelines to extract tutorial components from videos and transcripts to assemble mixed-media guides. Complementing these authoring tools, recent work demonstrates fully automated AR guidance from Large Language and Vision Models (LLVMs), enriching steps with embedded visual cues (Zhao et al., 2025). While this highlights the increasing integration of modern AI, its focus is on automated generation rather than end-user teaching and refinement.
Despite these advances, most systems target single-user, non-collaborative tasks. They seldom model role-dependent logic, joint timing, or the state of teammates and opponents. While systems like GhostAR (Cao et al., 2019) and VisCourt (Cheng et al., 2024) explore multi-agent choreography and tactic authoring in AR and MR, they primarily focus on configuring pre-defined interactions rather than learning an editable, closed-loop policy from narrated demonstrations.
2.2. Interactive Task Learning
In Interactive Task Learning (ITL), the goal is for a machine to learn task knowledge from a human teacher, often incrementally and through interactions that build an interpretable representation (Laird et al., 2017). Prominent HCI examples include ONYX (Ruoff et al., 2023), which enables users to teach natural-language interfaces by aligning speech and demonstrations with system capabilities, and VAL (Lawley and MacLellan, 2024), which uses LLMs in a neuro-symbolic pipeline to acquire editable, hierarchical procedures from conversation. These systems excel at software-centric or simulated tasks, emphasizing symbolically interpretable knowledge and conversational repair (e.g., clarifications, confirmations, undo). However, they typically do not ground learning in continuous physical trajectories or reason about multi-agent coordination and timing—challenges central to collaborative physical activities.
2.3. Interactive Program Synthesis
A parallel research thread focuses on synthesizing executable programs from end-user input, such as examples or demonstrations. Canonical work like FlashFill (Gulwani, 2011) demonstrated how carefully designed domain-specific languages (DSLs) could yield powerful transformations from a few examples. Subsequent HCI research has developed interaction models to make synthesized programs more understandable and editable. This includes techniques for disambiguating synthesized programs via navigation and active-example queries (Mayer et al., 2015), and augmented-example workflows that allow users to iteratively repair programs (Zhang et al., 2020).
Domain-specific systems show how end users can create and refine automations involving complex logic like loops and conditionals. Examples include Rousillon (Chasins et al., 2018) for hierarchical web scraping, DiLogics (Pu et al., 2023) for handling diverse input conditions, and MIWA (Chen et al., 2023) for mixed-initiative diagnosis and repair. In mobile computing, SUGILITE (Li et al., 2017) and PUMICE (Li et al., 2019) combine natural language with programming-by-demonstration (PbD) (Cypher, 1993) to synthesize procedures that users can inspect and modify. In data science, Wrex (Drosos et al., 2020) produces readable notebook code from examples, Falx (Wang et al., 2021) synthesizes visualization pipelines from example outputs, and recent interfaces like SQLucid (Tian et al., 2024) generate step-by-step, editable natural-language explanations to help non-experts understand and correct model-generated SQL.
3. Background
This work builds on recent advances in multimodal learning and probabilistic programming. In this section, we introduce the technical foundations that enable our system design.
3.1. Large Vision Language Models (LVLMs)
Large Vision Language Models extend large language models by processing visual inputs in addition to linguistic ones. LVLMs have demonstrated strong capabilities in aligning natural language descriptions with structured visual data, supporting tasks such as captioning, multimodal retrieval, and program synthesis (Radford et al., 2021; Yuan et al., 2021; Li et al., 2022). Their capacity to integrate symbolic and perceptual representations makes them well suited for interpreting narrated demonstrations, which combine visual demonstrations of physical activity with verbal explanations. In our work, LVLMs serve as the program synthesis tool.
3.2. Scenic Probabilistic Programming Language
Scenic is a probabilistic programming language designed for modeling complex physical environments with stochastic variation (Fremont et al., 2019, 2023)). By allowing users to specify distributions over 3-dimensional scene configurations and object behaviors, Scenic enables flexible yet interpretable representations of physical contexts. This stochasticity is particularly valuable for collaborative physical activities, where environmental conditions and agent interactions naturally vary across instances.
4. System Architecture
An overview of our system architecture is illustrated in Fig. 1. Throughout this section, we will use a running example where user teaches a soccer tactic on how to lure (or attract) an opponent defender away from the teammate to create space for the teammate to score.
Figure 2. Finite state machine representation of behaviors for our running example. The left figure shows a hierarchical FSM modeling a set of narrated demonstrations, and the right figure zooms into an hierarchical state of the FSM.:
4.1. Teaching Interface in Mixed Reality
Environment Modeling. To model and generate MR environments involving multiple entities with interactive behaviors, we use Scenic (Fremont et al., 2019), a probabilistic programming language designed to specify distributions of physical scenarios. The probabilistic aspect of the language will later be utilized to model variations in user’s narrated demonstrations. In this system, a Scenic program is provided, which models the initial scenes and the behaviors of environment agents, i.e. the entities that the user will interact with. [^1] It would be ideal if users can flexibly author the environments they desire from scratch. This authoring may be possible provided that user can teach just one’s own behavior, which we aim to examine through our user study. The full authoring is out of scope of this paper.
Action Recognition. Having large vision language models to understand sophisticated human actions, especially collaborative physical activities, is still an open challenge. Thus, we restrict the user actions to be triggered via MR controller buttons such that the system can easily recognize and log the user’s actions.
Embodied Narrated Demonstrations. User wears a MR headset which generates a pre-scripted environment overlaid on the physical world. The user is immersed in the environment physically interacting with environment agents. From this embodied perspective, the user provides physical demonstrations, interacting with environment agents, while verbally explaining the rationale behind actions. During the narrated demonstrations, the MR headset records the following data: (1) RGB video of the demonstrations recorded from virtual camera, (2) timestamped annotation of actions (e.g., [12:33] “User passed to Teammate”), (3) transcription of narration which consists of timestamped tokens (e.g., [12:34] “Therefore”, [12:35] “you”).
4.2. Program Modeling
4.2.1. Modeling Behavior as a Finite State Machine
The objective of program synthesis in our system is to model the user’s behavior from narrated demonstrations. We formally represent the behavior as a hierarchical, interrupt-driven finite state machine (FSM), a standard abstraction in program synthesis that compactly encodes contingencies via conditional branching. In our setting, states (nodes) denote actions, while directed edges denote preconditions—Boolean predicates over continuous variables (e.g. positions, orientations) of the physical world—that gate transitions. This abstraction allows the model to capture both discrete sequencing and coordinated behavior conditioned on the physical world state.
Returning to our running example, the FSM in Fig. 2(left) models two narrated demonstrations from the user teaching how to lure an opponent. The user’s narrates “So, in this scenario you need to lure the opponent to either sides of the field like this to receive a pass from your teammate. After you receive the ball and the opponent does come after you, then your teammate is open so pass to your teammate.” In the next narrated demonstration, user states “After you receive the ball, if the opponent does not budge and goes after your teammate, then you can shoot for the goal.” The FSM in Fig. 2(left) models the decision flow of the user in both of the narrated demonstrations as a FSM.
Hierarchical vs. Primitive States. There are two types of states: hierarchical and primitive. An hierarchical state represents a FSM which consists of primitive states that does not expand into its own FSM. In Fig. 2(left), we represent a behavior with hierarchical states whose transitions are modeled with the preconditions over the edges. Each hierarchical state represents a FSM consisting of primitive states and edges, which represents a micro-controller that executes the action represented by the hierarchical state. Fig. 2(right) zooms into MoveTo until t1 state in Fig. 2(left), where MoveTo is the primitive state that executes the action followed by an interrupt-driven and non-interrupt-driven edges. If the action MoveTo completes, then it transitions to the Wait state which idles and takes no other action until either preconditions p1 or p2 are satisfied. If a termination condition (t1) is satisfied while MoveTo action is in progress, then an interrupt occurs where the transition to Wait occurs regardless of whether MoveTo completes. This interrupt-driven transition models the contingent human behavior that reacts to dynamically unfolding physical environment. The FSM in Fig. 2 can be modeled as a Scenic program in Fig. 4.
Figure 3. Composing spatial constraints in the running example. Each panel shows a probability distribution over the field: (a) availability of a clear passing line, (b) positioning to the left of the defender, and (c) positioning to the right. These constraints are composed with logical operators (AND/OR) to produce the final distribution (d), which is sampled to determine where the agent should move to create space for a teammate’s pass as taught by the user.:
4.2.2. Library of APIs
The library of APIs defines the action space and models physical constraints, both of which are domain-specific and need to be tailored accordingly.
Constraint APIs for Logical and Spatial Reasoning. To logically and spatially reason about the state of the physical world, our library provides a set of constraints. We define a constraint as an API that returns either a boolean value or a spatial field over the workspace, indicating whether the current state of the world satisfies the constraint, or where in the workspace the constraint would hold.
For instance, a constraint DistanceTo(obj, ref, d, operator) returns True if object obj is at distance d from ref under the specified operator, or alternatively a probability distribution over the workspace indicating where obj’s location would satisfy the constraint. This constraint is useful in our running example for detecting when the opponent is applying pressure to the user—for instance, when the opponent is positioned too close.
Composition of Constraint APIs. Constraints, whether boolean or spatial, can also be composed with logical operators: AND, OR, and NOT. Boolean composition follows standard logic, returning a final True or False. For spatial fields, composition is defined as:
$A\land B=A\cdot B,\quad A\lor B=A+B,\quad\lnot A=1-A,$
where $A$ and $B$ are constraint spatial fields. The resulting distribution is a probability distribution discretized over the workspace, following the logical construction of the constraints. After normalization, this distribution can be sampled to produce concrete positions.
Returning to our running example, the user taught “So, in this scenario you need to lure the opponent to either sides of the field like this to receive a pass from your teammate.” Fig. 3 illustrates how spatial constraints can be logically composed to capture the conditions in our running example. To receive a pass from the teammate (blue), the user should not be standing behind the opponent (red). Thus, Fig. 3(a) assigns low probability (dark area) distribution to locations behind the opponent from the teammate’s perspective, while assigning high probability on all other locations (yellow). To move to either sides, the two constraints (Fig. 3(b) and (c)) model horizontal conditions assigning high probability distributions to either sides of the field. The gradation of colors reflect the Gaussian distribution which we encoded to these constraints. The composition of these constraints means applying the athematic shown above followed by a normalization as visualized in Fig. 3(d).
Action APIs
Action APIs define the action space of a user which can be either discrete (e.g. Pass) or continuous (e.g. MoveTo(destination)).
Figure 4. LVLM output Scenic program for running example. Italicized identifiers (precondition1, etc) are learned constraints.:
4.3. Program synthesis
4.3.1. Data Processing
We overlay visual cues onto the RGB videos of narrated demonstrations. These cues include a grid to visualize object positions relative to the workspace and each other, markers for user annotations (e.g., pointing to where a player should move to create space for a pass), object labels, and arrows to indicate orientation. Such overlays support the LVLM’s visual and physical reasoning. Videos are then subsampled as a sequence of images at a uniform rate according to the LVLM’s token context length.
The transcribed narration is grounded with annotations and logged events from the demonstrations, with additional text injected into the transcript to clarify verbal references, establish action order, and enrich the description. For instance, in our running example, the transcript becomes: “You should move to either side of the field [user marked coordinate $(x,y)$] to open up for a pass. [teammate passed to user] Then shoot if a clear line is available [user shot toward the goal]; otherwise, pass back to your teammate.” Grounding is achieved by aligning timestamps of word tokens with logged events.
4.3.2. Program Generation
The grounded multimodal streams—comprising the annotated transcript and videos with overlaid visual cues—are provided as input to the LVLM together with the library of APIs described in Section 4.2.2. A system prompt guides the model to synthesize a Scenic program that encodes the demonstrated behavior utilizing the available APIs. In doing so, the LVLM need to account for both variations and overlap across narrated demonstrations, producing a program that generalizes the user’s intent while adhering to the FSM-based modeling scheme.
Adding Temporally Aligned Narration to System Demonstration. Demonstrations alone can be ambiguous, as users may infer multiple, conflicting interpretations of the system’s intent. To address this, we introduce a Speak() API that allows the program to narrate its intent during execution. This narration occurs in real-time, synchronized with the corresponding actions, enhancing interpretability. Prior work in interactive program synthesis has explored natural language explanations through inline annotations (Angert et al., 2023) or step-by-step summaries (Liu et al., 2023a). Building on this, our system prompts the LVLM during synthesis to insert line-by-line rationales using Speak(text) calls, which encode the reasoning behind each local action. At runtime, each Speak statement briefly pauses the simulation, plays the narration via a text-to-speech model, and then resumes execution. Because these narrations are embedded directly within the program’s structure, they are temporally aligned with the behavior being demonstrated, enabling users to understand not just what the system is doing, but why it is doing it, in the moment. Returning to our running example, the synthesized program modeling the user behavior with narration is shown in Fig. 4(left).
4.4. User Inspection and Feedback to System
Our method allows users to amend the program through two primary code-free editing interfaces.
4.4.1. Decision Flow
Prior work shows that symbolic representations of program structure help non-programmers inspect logic and make targeted edits (Xie et al., 2024; Yen et al., 2024; Jiang et al., 2022; Kazemitabaar et al., 2024). We adopt this visualization approach by displaying the program’s FSM. For the program in Fig. 4, the user is shown the corresponding decision flow chart in Fig. 5(a). Nodes and edges include descriptions that users can reveal by clicking, which convey the rationale and underlying code logic without exposing the code itself. Users can annotate selected nodes and edges while providing verbal feedback, enabling direct edits to the program’s structure. The user might look at the decision flow chart in Fig. 5(a) and annotate the node for MoveTo and explain that ”You don’t need to move to the goal. Either make the shot or pass back.” Otherwise, if everything was fine with the decision flow chart, the user could just verbally state that ”no modifications are needed.”
4.4.2. Program Execution
The program runs as a Unity MR simulation, producing an embodied, narrated execution of the learned behavior. It controls an avatar—a proxy for the user—across variations of the environment encountered during teaching. As the behavior unfolds, the rationale and actions are narrated through the Speak actions defined in the code as seen in Fig. 4 (left). Users can pause at any time, point and click to annotate the environment, provide verbal feedback, and then resume—supporting multiple corrections within a single run. As in data processing for program synthesis, we capture both video and a grounded transcript of the user’s feedback, with interleaved annotations. Incorporating the Speak() outputs into the transcript further helps the LVLM localize feedback to specific points in the program logic, using the narration as a debug print statements. After looking at the program execution in Fig. 5(b) the user might pause at the frame shown and verbally explain that ”You should move further to the side of the field, around where I marked, to lure the opponent away [user annotated (x,y)(x,y)].”
4.4.3. Program Repair
We then provide the multimodal stream of feedback data including videos, annotations and grounded transcripts to the LVLM together with the candidate program to amend and the user’s initial narrated demonstration. The original prompt for program synthesis is augmented to instruct the LVLM on how to utilize feedback data and make localized edits.
Figure 5. User-in-the-loop feedback is supported through two complementary modes: (a) Decision Flow, which presents a finite state machine derived from the synthesized code, enabling structured inspection and annotation of program logic. Users are able to click and highlight edges and nodes by using ”Annotate.” When edges and nodes are highlighted yellow through annotation, our system will know which edgers and nodes need to be fixed. (b) System Execution, which replays the synthesized behavior under varied conditions with optional narrated Speak() actions, allowing users to validate fidelity in practice and provide situated corrections. Users can provide annotations in the environment by clicking on the field, leaving behind an X mark. Together, these modes balance interpretability and experiential validation, ensuring programs remain both logically consistent and behaviorally faithful.: :
Table 1. Research questions and measures.
Research Questions Metrics
RQ1: Do users find it intuitive to teach physical activities via narrated demonstrations with our embodied MR interface? 1. User Questionnaire
RQ2: Can users generate, on first attempt, an accurate program that models their behaviors? 1. Correctness score 2. Completeness score
RQ3: Can users efficiently improve the accuracy of the program via iteratively editing with our system? 1. Correctness score 2. Completeness score
RQ4: Do users find it intuitive to edit the program? 1. User Questionnaire
5. User Study
We aim to evaluate the efficacy of our system in supporting users with no programming background to accurately externalize their domain knowledge of collaborative physical activities as programs. In particular, we investigate the four research questions in Table 1.
5.1. Participants
We recruited 20 participants via email to our institution’s intramural, club, and varsity soccer groups. Inclusion required competitive soccer experience (at least competitive intramural, club, or varsity at the high-school or college level). We excluded individuals who only played recreational leagues, had a history of seizures, visual abnormalities, susceptibility to motion sickness or nausea, scalp skin irritation, or used wearable medical devices, consistent with the MR headset’s safety guidance. Each participant received a \$30 gift card.
5.2. Experiment Setup
Users used a Meta Quest 3 MR headset to teach soccer tactics at outdoor on a lawn (Fig. 1). When teaching soccer tactics in MR, users moved around and executed actions (e.g. pass, shoot, start/stop recording, pause/unpause scenario) using MR controllers. After users provided narrated demonstrations, the headset transmitted the collected data to the study personnel’s laptop. All program synthesis were conducted on the laptop. OpenAI 5.0-mini was used as LVLM. The decision flow diagrams and the program executions were also shown in the laptop to collect user feedback.
Participant ID Highest Soccer League Played Years of Coaching Experience Gender
1 Competitive 1 Male
2 Semi-Professional 0 Female
3 Competitive 1 Male
4 Competitive 0 Male
5 Competitive 0 Female
6 Semi-Professional 1 Male
7 Semi-Professional 2 Female
8 Semi-Professional 0.5 Male
9 Competitive 2 Female
10 Competitive 0.5 Female
11 Competitive 0.5 Male
12 Semi-Professional 7 Male
13 Competitive 0 Female
14 Semi-Professional 3 Male
15 Competitive 0 Male
16 Semi-Professional 3 Female
17 Competitive 1 Female
18 Semi-Professional 0 Male
19 Semi-Professional 3 Female
20 Semi-Professional 0 Male
Table 2. The summary of the participant backgrounds. The “competitive” league includes soccer club and competitive-level intramural leagues. The “semi-professional” league includes varsity or higher level. The “recreational” league includes recreational intramural leagues.
5.3. Study Design
We conducted a within subjects study where participants evaluated the system’s learning before and after they provided corrective feedback. We chose soccer as an example domain for complex collaborative physical activities as it involves highly dynamic coordination with multiple players. Three different soccer scenarios programs were prepared by the study personnels, where the behaviors of the environment agents (e.g. teammates, opponents) were pre-scripted. These three scenarios are derived and adapted from the England Football Learning (England Football Learning, 2025), a popular website for youth soccer coaches regarding soccer drills. Each participant embodied the coach avatar in MR to provide narrated demonstrations. The soccer scenarios were assigned to participants in a round-robin manner.
5.4. Study Procedure
Each participant completed a 60-minute one-on-one session. After watching a tutorial video (covering how to teach in AR, inspect, and correct system learning), participants had a hands-on tutorial in MR to familiarize the buttons on the controllers. Then, they watched a video informing them of their assigned scenario and the pre-scripted behaviors of the environment agents (e.g. teammates, opponents), and reviewed the rubric that would be used for evaluation of the learned programs later to ensure that their teaching covers the rubric. They then recorded two narrated demonstrations of the same tactical goal (re-recording allowed). The system synthesized a program and displayed a decision-flow diagram; if issues were identified, participants were allowed to provide at most one corrective-feedback and the LLM repaired the program. Next, the system showed three narrated demos of its learned behavior; if problems remained, a fourth demo was shown and participants were allowed to provide at most one feedback, prompting another repair. The restrictions on the number of user feedback was enforced due to limited study time. Participants evaluated the program pre-feedback and post-feedback using the rubric. After viewing both versions’ decision-flow diagrams, they sketched their ground-truth decision flow and compared it with the system’s baseline and post-repair diagrams. Sessions concluded with a brief user experience interview (refer to tutorial videos and rubrics in the Supplement).
Figure 6. The distributions of user scores related to each statement is shown. :
6. Results
6.1. Metrics for Quantifying the Alignment of Learned Program Behaviors with User Intent
We quantify the alignment in two dimensions. First, the correctness of the program behavior is evaluated
Correctness. This metric evaluates the program execution (rollouts). Users observe three narrated demonstrations of the program and scored them based on the rubric that we showed them before they provided their initial narrated demonstrations for teaching to ensure that the rubric is reflected in their teaching. Raw rubric scores were normalized to a percentage scale. The correctness score for each program execution $i$ is defined as
$C_{i}=\frac{s_{i}}{s_{\max}}\times 100,$
where $s_{i}$ is the participant’s rubric score and $s_{\max}$ is the maximum possible rubric score.
Completeness. This metric complements correctness. The program executions that users evaluated may not cover all branching code paths created by different conditional statements. Thus, completeness quantifies the alignment of the overall program structure (visualized as decision flow diagram) with the user intent. Recall, in the study, each user sketched their ground truth decision flow and compared it with the system’s 5.4. Let $G_{sys}=(V_{sys},E_{sys})$ denote the set of nodes and edges in the system-generated decision flow, and $G_{gt}=(V_{gt},E_{gt})$ denote the ground-truth’s. Completeness is calculated as the percentage of ground-truth elements (nodes and edges) that also appear in the system’s flow:
$\text{Completeness}=\frac{|V_{sys}\cap V_{gt}|+|E_{sys}\cap E_{gt}|}{|V_{gt}|+|E_{gt}|}\times 100.$
6.2. Statistical Analysis Method
Wilcoxon signed-rank test (Woolson, 2007) was used to evaluate to conduct one-sided hypothesis test on whether alignment improved post-feedback within subjects. Wilcoxon test was used instead of paired t-test because the resulting distribution of metric values were non-Gaussian. The scipy.stats.wilcoxon function from the SciPy library was used.
6.3. User Experience on Teaching via Embodied Narrated Demonstrations (RQ1)
95% of the users agreed that it was easy to teach soccer tactics via narrated demonstrations (Fig. 6). However, users also pointed out some limitations of the teaching interface, where two themes emerged. First, 8 out of 20 participants raised issues regarding the pre-scripted environment agents’ behaviors not aligning with their expectations to provide realistic teaching. The pointed out unrealistic movement (linear/stiff paths, stationary or slow defenders) and timing/latency that diverged from typical live play. For example, one noted “Players can only run in linear paths… addition of curved paths would be more realistic” (P2). Another shared “I didn’t feel that I was actually under pressure… unrealistic how slow [defenders are] moving” (P15). The other theme related to the lack of real-time feedback from the system to the participants while they were teaching. A few participants wanted two‑way feedback/acknowledgment from avatars during the teaching phase. “There’s no feedback from the [VR] players… felt a little off compared to real life.”(P13) “It felt a little bit like I was talking to the void… [I wanted] some way of having… feedback.” (P18)
Figure 7. Box plots showing the average and standard deviation of percentage score of the initial synthesis before feedback, and the percentage score of the synthesis after feedback. A score of 100% on correctness means that the system execution performed all the tasks correctly according to the user, scored from a pre-determined rubric for that scenario. A score of 100% on completeness means that the decision flow chart outputted by the system is 100% aligned with the user’s own decision flow chart.:
6.4. Initial System Learning from Narrated Demonstrations of Users (RQ2)
Pre-feedback alignment of learned programs is shown in Fig.7 (top left) and Fig.7 (bottom left). The wide variance likely reflects differences in instructional quality. Among the 10 participants whose programs achieved correctness or completeness scores greater than 80%, their narrations shared two traits: (i) 9/10 explicitly stated the purpose behind actions (e.g., Now your teammate will make an overlapping run, and you want to pass it to them so that they can score the goal.” (P5)), and (ii) 7/10 articulated multiple possible outcomes to keep executions correct under varied conditions (e.g., We’re gonna see if the shot is open, I’m gonna take it, but if it’s not, now I’m gonna pass it. Uh, or, let’s say I got covered, I could dribble a bit, and then I can take the shot.” (P15)). In contrast, 7 participants whose learned program scored less than 50% in either metrics exhibited the opposite, either did not explain the rationale (3/7) or did not consider multiple outcomes (4/7).
6.5. System Learning from Corrective Feedback of Users (RQ3)
[…]
6 of 20 participants did not improve on either correctness or completeness after feedback. These outcomes stemmed from three sources: (1) Insufficient APIs: 2 out of 7 participants tried to correct the trajectory curvature of movements taken by the program-controlled avatar. However, our APIs supported only optimal path planning/control and, therefore, could not accommodate this feedback. (2) Context loss during repair: To reduce synthesis time, we omitted original narrated videos from the LLM repair prompt for feedback on system’s narrated demonstrations. Two of 20 participants then gave feedback that seemingly contradicted their earlier teaching (assuming the system retained that context), leading the LLM to delete intended branches. For example, a unique issue occurred for P17, who ended feedback with “Okay, we can stop here,” which the LLM took literally, producing a truncated program. This accounted for 2/7 non-improving cases. (3) Deadlocks: In 4 of 7 participants, the synthesized program deadlocked because a learned condition was never satisfied. Our system did not explain which condition blocked execution, preventing effective debugging by the users.
Figure 8. Mean Likert ratings (1–7; higher = more agreement) before vs. after feedback. Left: the avatar’s decision flow accurately reflected what I intended (Before: 4.7, After: 5.4). Right: both the avatar’s decision flow and demonstrations, were accurate enough to teach someone else (Before: 3.5, After: 4.6). One participant did not provide feedback and was excluded.:
6.6. User Experience on Program Editing (RQ4)
While correcting system learning, 16/20 participants found the decision flow easy to understand and debug: “Felt like… if-then… I liked that the decision flow was… logically laid out.” (P7) “Correcting the decision flow was pretty smooth… not challenging.” (P14) However, 4/20 reported difficulty with comprehension, noting confusion distinguishing arrows, edges, and nodes: “Understanding what the arrows were versus the… edges versus the nodes was a little confusing in the beginning” (P5).
All 20/20 participants found the system’s narrated demonstrations easy to follow and diagnose in general: “Very intuitive to make annotations and identify where there were errors” (P1). “It was easy to understand what the avatar got from me based on the narrations and demonstrations” (P8). However, one participant cited unclear terminology was used in one of the system narrations: “Language… ‘recycle the attack’… a little bit unclear” (P18). Regarding UI, while 17/20 participants reported that they were able to sufficiently express their feedback, 3 participants requested richer feedback tools for specifying trajectories: “It would help to be able to draw trajectories… more sophisticated motion” (P10).
7. Discussion
[…]
While the flexibility of free-form narrated demonstrations could allow users to teach their unique strategies, we observed that this freedom could also introduce challenges in externalizing tacit physical knowledge. The initial program synthesis exhibited wide variability, with standard deviations of 24.02% for correctness and 28.85% for completeness. Our analysis suggests this variability stems from differences in teaching styles; participants who explicitly articulated the rationale behind their actions and described potential contingencies tended to produce programs that were more aligned with their intent. Conversely, programs synthesized from narrated demonstrations that omitted this contextual information were tend to perform poorly.
In future work, this omitted information could be elicited by incorporating a framework inspired by Theory of Mind (Premack and Woodruff, 1978; Baker et al., 2017; Rabinowitz et al., 2018). Instead of passively receiving instruction, the system could actively reason about the user’s mental model of the task. For instance, an LLM could analyze the narrated demonstrations to identify points of ambiguity or underspecified conditions. It could then generate targeted questions to probe the user’s intent, such as, “I noticed you moved left. Was that to create space, or were you reacting to your teammate’s position?” By modeling the user’s knowledge and identifying gaps, the system could guide the user to provide the contingencies and rationales needed for robust program synthesis, shifting the interaction from a one-way demonstration to a collaborative dialogue.
Our results demonstrate that participants were able to successfully interpret and correct the synthesized programs to better align them with their intent. This corrective feedback loop proved effective; even with a single round of feedback—which 16 of 20 participants provided due to time constraints—program completeness increased significantly by 16.68% (SD=18.52, p=0.03). This quantitative improvement was echoed in participants’ subjective ratings of the decision-flow structure, which also improved post-feedback (Fig. 8). Participants were consistently able to steer the synthesized program toward their intended high-level logic for the soccer tactics (Fig. 7, right). In interviews, several participants noted that additional, iterative rounds of feedback—rather than a single round—would likely yield further improvements.
However, the study also revealed key limitations. Unlike completeness outcomes, the improvement in program correctness showed upward trend of 6.87% (SD=22.42) on average but was not statistically significant (p=0.10). Our analysis pointed to two fundamental issues. First, the expressiveness of the underlying API library was a significant constraint. In some cases, participants were unable to improve program alignment because the provided APIs lacked the necessary functions to implement their intended corrections. Recall that our system restricted LVLM to not arbitrarily implement APIs and only use the ones provided. While this restriction may be helpful in reducing potential hallucinations of LVLM, it could also constrain the expressivity in modeling user behavior in trade-off. This suggests that for a system to be broadly applicable within a domain, its API library must be iteratively developed through an in-depth analysis of user needs.
Another major limitation of the system was its failure to explain program deadlocks. On 4 out of 6 occasions where users fail to better align the programs with their intent, users experienced deadlocks where the programs halted in the middle because one of its preconditions are not satisfied. The system could only narrate the program logic and lacked a mechanism to detect and communicate which specific condition(s) were causing the halt. This made it particularly challenging for users to debug when preconditions involve multiple constraints. This limitation could be mitigated by augmenting the synthesis prompt to generate routines specifically for detecting deadlocks and communicating the source of the error to the user.
While the majority of participants (95%) found the narrated demonstration interface easy to use, their feedback also highlighted a key question regarding the necessary scope of instruction for collaborative tasks. Our study utilized environment agents with pre-scripted behaviors, which several participants found constraining when the agents’ actions diverged from their tactical intent. This suggests that for complex collaborative activities, modeling the user’s behavior in isolation may be insufficient for fully externalizing their knowledge. A natural extension of our work would be to allow users to teach or modify the behavior of other agents in the environment. While our system’s architecture could be adapted for this purpose, such an extension would introduce additional system complexity and potentially increase the user’s cognitive and physical load, requiring careful investigation into mitigation strategies.
While the majority of participants (90%; 18/20) found the process of correcting the model to be straightforward, a key challenge emerged in understanding what the model had learned, which in turn complicated debugging. Five participants reported that the decision-flow diagrams became difficult to interpret as they grew in complexity, often containing redundant nodes and edges. This issue likely stems from the absence of a parsimony objective during program synthesis and visualization. Future iterations could improve interpretability by imposing minimality constraints on the program structure and its visual representation within the LLM prompt. Furthermore, while most participants inferred behavior from the system-generated narrated demonstrations, some were confused by technical jargon (e.g., “check for a path width of 2 meters”). This highlights the need to extend the interface to allow users to inquire about and clarify ambiguous terms, making the system’s communication more transparent.
8. Conclusion
In this paper, we presented an interactive program synthesis system that enables non-programmers to teach complex, collaborative physical tasks. We framed this problem as one of program synthesis, where the system learns an editable, interpretable program that represents the collaborative activity. Our core contribution is an interaction model that uses narrated demonstrations as a unified, bidirectional modality: users teach and correct the system through paired physical actions and language, and the system communicates its learning using the same modality, removing the need for users to interact with raw code. Our study showed feasibility of this approach. Users successfully taught multi-agent soccer tactics and significantly refined the synthesized programs to align with their intent. Our findings have implications for the design of future interactive systems for task learning, highlighting the need for mechanisms to more actively elicit tacit knowledge from users and to expand the instructional scope to include the behavior of collaborating agents. This work contributes a novel interaction framework and demonstrates the feasibility of synthesizing executable models of complex human collaboration directly from end-user instruction.
Appendix A User Study Supplement
A.1. Tutorial Video
Here is the link to the tutorial video that all participants watched at the beginning of the study. This video covers how to teach, inspect, and correct the system via narrated demonstrations.
A.2. User Scenarios
In the study, each participant was assigned one of three pre-scripted scenarios. For each scenario, we prepared a tutorial video explaining the behaviors of the teammate(s) and opponent(s) as well as the tactical goals to be achieved. The shared drive contains the three tutorial videos. The scenarios were: (1) lure, (2) overlap, and (3) distribute.
A.3. Rubric for Each Scenario
Here is the documentation of the three rubrics for the corresponding scenarios, which participants used to score the correctness of learned programs after watching their executions. Each participant was shown the rubric for their assigned scenario before teaching the system to ensure instruction that could yield full scores. Each rubric consists of multiple statements, each rated 0, 1, or 2: a score of 2 indicated the program completed the statement exactly as intended, 1 indicated partial or approximate completion, and 0 indicated failure. Participants watched three program executions and scored each with the rubric; their reported score was the average of the three. These results are shown in Fig. 7 (left two plots).
A.4. User Interview Questionnaire
  1. (1)
    How was your experience teaching in augmented reality?
    1. (a)
      What did you find intuitive?
    2. (b)
      What did you find challenging?
    3. (c)
      How can we improve the system?
  2. (2)
    How was your experience correcting the avatar’s behavior by commenting on the avatar’s decision flow?
    1. (a)
      What did you find intuitive?
      1. (i)
        Was it easy to understand the decision flow?
    2. (b)
      What did you find challenging?
    3. (c)
      How can we improve the system?
  3. (3)
    How was your experience correcting the avatar’s behavior by directly annotating and commenting on the avatar’s narrations and demonstrations on the soccer field?
    1. (a)
      What did you find intuitive?
      1. (i)
        Was it easy to understand what the avatar learned based on its narrations and demonstrations?
    2. (b)
      What did you find challenging?
    3. (c)
      How can we improve the system?
  4. (4)
    How was the avatar’s learning compared to what you intended to teach?
    1. (a)
      How would you think of the avatar teaching someone else based on what it learned from you?
      1. (i)
        What needs to be improved?
Figure 9. Three plots overlay participants’ demonstration trajectories with execution trajectories from the learned probabilistic programs, each corresponding to one of the three user scenarios in Appendix A.2.
A.5. Generalizability of Learned Programs
The learned Scenic probabilistic programs did not simply replay participants’ narrated demonstrations. In the study, each participant provided two demonstrations of the same task, and these variations were modeled as distributions in Scenic programs. Semantically, a Scenic execution samples a concrete scenario from the distribution it encodes and generates it in simulation. Fig. 9 overlays two demonstration trajectories of a participant with five program execution trajectories for each of the three scenarios: overlap (left), lure (middle), and distribute (right), corresponding to Participants 9, 10, and 11. These participants produced learned programs that achieved 100% correctness and completeness via interactive program synthesis.
A.6. System Implementation
A.6.1. Teaching, Inspection, Editing Interfaces
The details of our three interfaces are explained in our tutorial video linked in Appendix A.1.
A.6.2. APIs
Our study personnels consisted of soccer players with the history of playing in the competitive soccer leagues. We used our domain knowledge of soccer to model the sets of action space and the physical constraints as APIs. For action space, we provided APIs that execute moving/dribbling, passing, shooting a goal, and triggering a teammate to pass the ball to the user. For modeling the physical constraints, we APIs that can validate conditions such as (1) ball possession, (2) distance relations, (3) horizontal/vertical relations, and (4) existence of passing lane between two players of interest. To validate these conditions, the APIs were internally given access to the world states (e.g. positions, orientations, ball possession of each player) in mixed reality simulation at each timestep. The annotations of all APIs were provided to the large vision language model (LVLM) for program synthesis. For example, Fig. 10 shows the annotation of DistanceTo() API which checks for the distance relation between two players of interest.
A.6.3. System Prompt
The system prompt used to instruct LVLM generate programs for our user study is shown in Fig. 11.
Appendix B Applicability of Our System to Other Collaborative Physical Domains
Our system architecture could support other collaborative physical domains where agents must coordinate in space and time, because it decouples domain-agnostic synthesis/inspection/repair loops from domain-specific action and constraint primitives. Narrated demonstrations are still used in the teaching interface; the LVLM compiles multimodal traces into Scenic programs whose stochastic scene specifications capture variation in strategies and environments. Users then inspect and correct behaviors via decision-flow visualization and narrated executions. Retargeting to another domain entails (i) modeling the target environment and agents in Scenic and (ii) replacing the API library.
Example in Manufacturing Domain
We keep the synthesis, inspection, and repair loop unchanged and swap the API library for this domain with three actions and four constraints. In this scenario, the user provides teaching for a humanoid robot as to how to collaborate with a human worker.
Action APIs: goTo(location), pick(item, from),
swapBuckets(new_bucket, old_bucket, station).
Constraint APIs:
BelowThreshold(bucket, n):
Trigger predicate from vision/weight sensing; true when the measured part count in bucket is <n<n.
HumanReady(worker):
returns true when the worker signals it is safe to perform the swap.
Instruction from the narrated demo (“when the human worker’s bucket containing assembly parts has less than 10 parts, go fetch a new bucket full of the assembly parts from the supply station and swap the human worker’s bucket under permission”):
behavior UserBehavior():
while True:
  if BelowThreshold(worker_bucket, 10):
    do Speak(‘‘the worker’s bucket is running low on assembly parts. Fetch another bucket from the supply station’’)
do goTo(SupplyStation)
do Speak(‘‘pick up another bucket full of parts’’)
do pick(full_bucket, SupplyStation)
    do Speak(‘‘Return to the worker’s station.’’)
do goTo(WorkerStation)
    do Speak(‘‘Wait until the worker permits the bucket swap.’’)
do Wait() until HumanReady(worker)
    do Speak(‘‘The worker gave permission. Swap the bucket.’’)
do swapBuckets(full_bucket, worker_bucket, WorkerStation)
Figure 10. An annotation of an API we provided to a large vision language model.
Figure 11. The system prompt used in the user study for instructing LVLM how to code.